Would someone please help me solve this problem:
y-xy=-10
x-2y=-7
Thank you.

Answer by stanbon(48558)

1st: y-xy=-10
2nd: x-2y=-7
---------
Solve 2nd for x to get: x=2y-7
Substitute into 1st to get:
y-(2y-7)y=-10
y-2y+7=-10
-y=-17
y=17
Substitute into x=2y-7 to solve for x.
x=2(17)-7
x=27
SOLUTION:
x=27; y=17
Cheers,
Stan H.